DR MENTO HAS A CLASS OF 80 STUDENTS. FOR A GROUP PROJECT SHE WANTS TO DIVIDE THE STUDENTS INTO GROUPS OF 6, 8, 10. IS THIS POSSIBLE? EXPLAIN YOUR ANSWER.
Answer by vleith(2983)

No
Let's factor 80
2*40 = 2 * 2 * 20 = 2* 2 * 2 * 10 = 2 * 2 * 2 * 2 * 5
80 can only be 'grouped' into groups that are products of the prime factor above.
10 = 2 * 5 so that works
8 = 2 * 2 *2 so that works too
6 = 2 * 3 but 3 is not a factor of 80, so groups of 6 won't work